OpenStudy (anonymous):

The lengh of a rectangle is fixed at 23 cm. What widths will make the perimeter greater than 96 cm ? The width must be greater than ____ cm

OpenStudy (saifoo.khan):

This one... hmmm.. Let me show you.

OpenStudy (saifoo.khan):

This length of both sides = 23*2 = 46 Now, the width of both sides must be: 96-46 = 50 So the width of one side = 50/2 = 25

OpenStudy (saifoo.khan):

Check, 2(25 * 23) = 96.

OpenStudy (saifoo.khan):

Now the width must be greater than 25..

OpenStudy (saifoo.khan):

are you getting it?

OpenStudy (anonymous):

yeah now iam

OpenStudy (anonymous):

but im still getting the wrong answer

OpenStudy (anonymous):

show me the ending

OpenStudy (saifoo.khan):

Oh, okay, let me type.

OpenStudy (saifoo.khan):

WAIT. that's the ending. LOL.

OpenStudy (saifoo.khan):

25 is the final answer.

OpenStudy (anonymous):

oh ok i had 28 i was close

OpenStudy (saifoo.khan):

they asked the minimum value of it..

OpenStudy (saifoo.khan):

Cool. Do u want more help? i can help you through!!! :)

OpenStudy (anonymous):

yes hold up

OpenStudy (saifoo.khan):

:D

OpenStudy (anonymous):

for this problem 7(w-2)=49 i got the answer w=0

OpenStudy (saifoo.khan):

Nope, let me show u a shortcut to solve.

OpenStudy (saifoo.khan):

7(w-2) = 49 \[(w-2) = \frac{49}{7}\]\[w-2 = 7\]\[w = 7+2\]\[w=9\]

OpenStudy (anonymous):

tthat helped me a lot i was doing it wrong

OpenStudy (anonymous):

i have a couple more

OpenStudy (saifoo.khan):

Sure, i can help!!

OpenStudy (saifoo.khan):

You can also post a screenshot or photos of your questions here.

OpenStudy (anonymous):

ok cool

OpenStudy (saifoo.khan):

Thank God you are back.

OpenStudy (anonymous):

in 1993 the life expectancy in males in a c ertain country was 67.3 years, in 2000 it was 69.9 years , let e represent the life expectancy in year t and let t represent the number of years since 1993 . The linear function e (T) THAT FITS THE DATA IS ? tHIS WAS MY ANSWER E (t)= 0.7t+67.3 E(16)=68

OpenStudy (saifoo.khan):

Dang! that is one of the 2 questions which are hard as diamond. :O

OpenStudy (saifoo.khan):

Let me See.

OpenStudy (saifoo.khan):

I guess you are correct.

OpenStudy (saifoo.khan):

Do you have more?

OpenStudy (anonymous):

yup

OpenStudy (saifoo.khan):

:D Post them!!! XD

OpenStudy (anonymous):

solve using the multiplication principle -6x>1/13 i got the answer - 1/78

OpenStudy (saifoo.khan):

\[-6x > \frac{1}{13}\] \[x < -\frac{1}{78}\]

OpenStudy (saifoo.khan):

Stay careful about the inequality sign.

OpenStudy (anonymous):

ok cool i have more hold up

OpenStudy (anonymous):

graph the equation using the slope and the y-intercept y=10/3x+3

OpenStudy (saifoo.khan):

http://www.twiddla.com/639825

OpenStudy (saifoo.khan):

Come here.^

OpenStudy (anonymous):

p+q/8 for p=44 and q =4 i got 76

OpenStudy (saifoo.khan):

\[\frac{p+q}{8}\]Like that?

OpenStudy (anonymous):

yup

OpenStudy (saifoo.khan):

Alright,

OpenStudy (saifoo.khan):

You got it wrong! :( sorry.

OpenStudy (saifoo.khan):

Answer is 6

OpenStudy (anonymous):

how did u get that

OpenStudy (saifoo.khan):

\[\frac{44+4}{8} = \frac{48}{8} = 6\]

OpenStudy (anonymous):

i have more

OpenStudy (saifoo.khan):

lol, im doing all your mathlab H.w. No problem Im continuing helping!! :)

OpenStudy (saifoo.khan):

where are your questions?

OpenStudy (anonymous):

determine whether (-1,-7) is a solution of 4x-4y=6

OpenStudy (saifoo.khan):

Insert the values in the equation, 4x - 4y = 6 4(-1) - 4(-7) = 6 -4 + 28 = 6 24 = 6 As we are getting result which is not equal to 24, then this is Not the solution of the equation.

OpenStudy (anonymous):

i got the same answer hold up i have more

OpenStudy (anonymous):

whats mathlab

OpenStudy (saifoo.khan):

C'mon are u kidding? LOL.

OpenStudy (anonymous):

serious at my school its mathzone

OpenStudy (anonymous):

u ready for the other questions

OpenStudy (saifoo.khan):

what school are you in?

OpenStudy (saifoo.khan):

Yes, im ready.

OpenStudy (anonymous):

drexel

OpenStudy (anonymous):

freshmen

OpenStudy (saifoo.khan):

how can you have same questions?

OpenStudy (saifoo.khan):

well, i help many online college students.. they have the same questions as yours!!

OpenStudy (anonymous):

wow the same exact

OpenStudy (saifoo.khan):

Yes, the same. and i have done these many times!

OpenStudy (anonymous):

here is my next questi

OpenStudy (anonymous):

find the slope and the y -intercept f(x)= -10x-8 my answers are the slope -4/5 y- (0,-8)

OpenStudy (saifoo.khan):

Nope, slope is -10. y = (0,-8)

OpenStudy (anonymous):

ok i have more hold up

OpenStudy (saifoo.khan):

How many do u have man?

OpenStudy (saifoo.khan):

lol, just wondering...

OpenStudy (anonymous):

graph the line containing the pair of points and find the slopes (0,5) (-4, -2) i have m=-4 i didnot graph yet

OpenStudy (saifoo.khan):